Suppose that the United States and Canada can each produce two products: lumber and beef. Create a table like the one below, showing labor requirements per unit of output for each country. (Hint: Choose numbers for each country that are easily divisible by one another.)

Labor Requirements per Unit of Output

United States Canada
Lumber
Beef

What does absolute advantage mean? How do you calculate absolute advantage? In what output(s) does the U.S. have an absolute advantage? Explain using the data from your table. In what output(s) does Canada have an absolute advantage? Explain using the data from your table.

What does comparative advantage mean? How do you calculate comparative advantage? In what output(s) does the U.S. have a comparative advantage? Explain using the data from your table. In what output(s) does Canada have a comparative advantage? Explain using the data from your table.

What product should each country export? Why?
Answer #1
Country Lumber Opportunity cost in terms of Beef Beef opportunity cost in terms of Lumber
US 4 1 lumber = 8/4 =2 labour for beef 8 1 beef = 4/8 = 0.2 labour for beef
Canada 6 1 lumber = 3/6 = 0.2 labour for beef 3 1 beef = 6/3 = 2 labour for lumber

Absolute advantage means the country which can produce maximum quantity using same amount resources than other country will have absolute advantage in production of that particular good.

US has an absolute advantage in production of 1 unit of Lumber because they can produce one unit using 4 labour

Canada has absolute advantage in production of 1 unit of beef because they can produce it with 3 units of labour

Country which can produce a good at low opportunity cost is having comparative advantage in production of that good

In this case Canada has comparative advantage in production of Lumber and US in beef.

Country which have comparative advantage in that good should specialize in production of that good and trade with other country for another good.

US should export Beef as they have comparative advantage in production of beef and similarly Canada should export Lumber.